Guidelines for PET Technical Reports

C. Wayne Mastin^{∗}

February 23, 2000

1 Introduction

Your report should have the title and author on the first page, immediately followed by the text. This page can be used as an example. We will prepare a separate cover page and an acknowledgement/disclaimer page. The standard acknowledgement will read:

Work funded by the DoD High Performance Computing Modernization Program ERDC Major Shared Resource Center through

Programming Environment and Training (PET)

Supported by Contract Number: DAHC 94-96-C0002

Computer Sciences Corporation

If the work is jointly sponsored, please provide the necessary information (e.g. agency and contract no.) and we will include it in the acknowledgements. The disclaimer will read:

Views, opinions, and/or findings contained in this report are those of the author(s) and should not be construed as an official Department of Defense position, policy, or decision unless so designated by other official documentation.

The acknowledgement and disclaimer texts are provided for your information. We will print this page and include it in all reports. You should include a separate acknowledgement of computer resource support if you used HPCMP computer systems in your work. The text for that acknowledgement should appear after the last section and before the references. See the example toward the end of this document.

The author’s affiliation may be put in a footnote as illustrated here or placed after the author’s name. You may include an abstract if you wish, and that should be placed between the author’s name and the first section of the manuscript.

2 Manuscript

Your report may be prepared using HTML, Latex, or about any word processor. We need the electronic version of your manuscript. You may send it to us by email or make it available on a FTP or Web site. If you are using HTML, make sure you include any GIF files so we can reproduce the figures. If you are using Latex, you will need to send us any style files you use as well as the EPS files for the figures. To insure that we have the manuscript laid out the way you wish, you may want to also send us a hard copy of the manuscript.

∗PET On-Site Academic Lead, Computer Sciences Corporation

Figures and tables may be included in the body of the manuscript or collected together at the end. The final section should be followed by the acknowledgement of computer support and a list of references, if appropriate. If you choose to put figures at the end of the paper, the references should precede the figures. The references should also precede any appendices.

The manuscript should be single-spaced and the font should be of a size normally found in a technical report or journal article, about 10 to 12 pt. A one-inch margin should be allowed on all sides.

A sample report on PET training activities prepared by Mastin [?] is available as an example. The report is available either as a Latex text file or a Postscript file. Requests may be addressed to Wayne Mastin, mastin@wes.hpc.mil. All PET technical reports are available on the ERDC MSRC web page at http://www.wes.hpc.mil/pet/tech reports/available.htm.

Acknowledgement

This work was supported in part by a grant of HPC time from the DoD HPC Modernization Program.

References

[1] C. W. Mastin, “1999 ERDC MSRC PET Training Activities,” ERDC MSRC PET TR 00-05, Vicksburg, MS, January 2000.